OpenStudy (anonymous):

Write y = 4/5x+ 8 in standard form using integers.

OpenStudy (anonymous):

please help

OpenStudy (dtan5457):

Integers are whole numbers. And you have a 4/5, so you'll need to multiply everything by 5 to get rid of the fraction.

OpenStudy (dtan5457):

y=4/5x+8 5y=4x+40 Now get that in standard form which is ax+by=c

OpenStudy (dtan5457):

Got it from here?

OpenStudy (anonymous):

im confused because you said ax+by=c but you put 5y=4x+40 I don't understand

OpenStudy (dtan5457):

In standard form of ax+by=c All it really means is to move x and y to one side and leave the whole number on the other. 5y=4x+40 is NOT in standard form yet

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so 4x+5y=40 is correct

OpenStudy (dtan5457):

Why is the 4x positive?

OpenStudy (anonymous):

its suppose to be positive

OpenStudy (dtan5457):

5y=4x+40 If you subtract the 4x on both sides I don't see why it's positive

OpenStudy (anonymous):

4x + 5y = 40

OpenStudy (anonymous):

I mean -4x+5y=40

OpenStudy (dtan5457):

I'm going with 5y-4x=40

OpenStudy (dtan5457):

Yours work too, same thing.

OpenStudy (anonymous):

ok thank you so much